OpenStudy (anonymous):

Find the area of a circle with an area of pi (9x^2+48x+64)

OpenStudy (anonymous):

I can't tell you what the area is unless you tell me what x is..but I CAN tell you that the radius can be expressed as 3x + 8 since area = pi*r^2 = pi*(9x^2+48x+64) = pi*(3x + 8)(3x + 8) = pi * (3x + 8)^2

OpenStudy (anonymous):

It doesnt give me x

OpenStudy (anonymous):

Then you can't answer the question.

OpenStudy (anonymous):

But 3x +8 is an answer choice

OpenStudy (anonymous):

That would be the answer to "What is the radius?" not "What is the area?"

OpenStudy (anonymous):

Maybe there was an error idk.

OpenStudy (anonymous):

Thanks anyways

OpenStudy (anonymous):

could be...good luck!